OpenStudy (anonymous):

How many grams of sugar are in six dozen doughnuts if each doughnut contains 75 g of sugar?

OpenStudy (anonymous):

5400?

OpenStudy (anonymous):

A dozen of anything is 12, so 6 dozen = (6*12) = 72 doughnuts. If one doughnut is 75g of sugar, 6 dozen = 72(75g sugar)

OpenStudy (anonymous):

So yes, 5400g sugar

OpenStudy (anonymous):

thanks :)